# Wave pulses with unusual asymptotical behavior at infinity

### v1

@@ II Hemispherical unidirectional pulse
We calculated the Poynting vector and energy density of $U\left(\rho,z,t\right)$ according to known formulas (see, e.g., Eqs. (1)-(5) of Ref. [18]) for scalar fields, as well as the total energy. The latter is finite. The former two have normal asymptotics in the sense of finite values of the limit of the type in Eq. (2) as they have to be since both consist of derivatives of $U\left(\rho,z,t\right)$. Despite the unidirectionality of $U\left(\rho,z,t\right)$—integration as a linear operation preserves the unidirectional spectrum of $u\left(\rho,z,t\right)$— in certain spatial regions the energy flux turned out to be directed opposite to the azis $z$, and at some points the velocity of such an energy backflow even reached almost the value $-c$.
@@ III Fractional splash pulse
Eqs. (15)-(17) state that if $t\rightarrow+\infty$ in other direction than along the positive axis $z$, the pulse exhibits normal asymptotics. Even if one modifies Eq. (17) by replacements $\rho=ct/\sqrt{2}+\Delta,\,z=ct/\sqrt{2}+\Delta$ in order to keep the point (running along the diagonal direction) close to the pulse’s top, the expression for the limit remains finite.
@@ V Discussion and study of physical feasibility
The following results were obtained : for all possible choices of $\vec{m}$, excluding $(0,0,1)$, the $x$-components and $y$-components of electric and magnetic fields both have abnormal asymptotics when $z\rightarrow\infty$ while the $z$-components do not. Hence, $E_{x}$, $E_{y}$, $B_{x}$, and $B_{y}$ behave similarly to $f\left(\rho,z,t\right)$, see Eqs. (14), (15), (28), and (29). In other words, when propagating in the positive direction of the axis $z$ they decay as $\sim 1/(z=ct)^{3/4}$, i.e., slower than the normal decay $\sim 1/(z=ct)$. If instead of the power $3/4$ another value between $1/2$ and $1$ is used in the definition of the splash pulse, the law of the decay changes correspondingly. Fig. 7 as an example of behavior of $E_{x}$ shows that in comparison with the scalar-valued wave function in Fig. 3 the bipolar profile has become richer. This is due to the circumstance that according to Eq. (30) the electromagnetic fields are expressed from $f\left(x,y,z,t\right)$ through double temporal and spatial derivatives. It is interesting to note that only $\partial^{2}_{z,z}$ and $\partial^{2}_{t,z}$ lead to the terms with abnormal asymptotics. In Fig. 7 there is a negative peak at a location $z<0$, which is stronger than the main positive peak. There may arise a doubt as to how then the expanding pulse’s edge at the domain $z<0$ decays faster than that in the domain $z>0$. However, a version of the plot in Fig. 6 that we made for $ctE_{x}\left(0,0,z,t\right)$ demonstrates that with growing of the value of $z>0$ the peak strength increases indefinitely while its counterpart at $z<0$ does not, i.e., both behave in accordance with Eqs. (14) and (15).
@@
The fact that the fractional splash pulse is not unidirectional does not imply that its physical feasibility is impossible. Its backward-propagating components can be suppressed through appropriate adjustments of the parameters $a_{1}$ and $a_{2}$. A practical approach to addressing this issue is to generate finite-aperture approximations of the pulse, similar to what was done in the experimental realization of the focus wave mode [28, 29].
As to the primitives in Eqs. (3) and (9), the electromagnetic and acoustic physical fields derived from them possess normal asymptotics. Therefore these primitives are mainly of theoretical interest.
@@ VI Conclusion
The solutions to the scalar wave equation with unusual asymptotic behavior in the far zone considered in this article are clearly of interest in mathematical physics. It is remarkable that there is a number of solutions which decay with propagation distance slower than the common inverse proportional law and yet have finite energy. Especially interesting is the family of fractional splash pulses. Indeed, it can be easily verified that if one changes in Eq. (13) the power $3/4$ with any other value between $1/2$ and 1, the pulse has slower than normal decay and at the same time its wave function is square-integrable and has finite total energy. These characteristics are passed on to physical electromagnetic as well as to acoustic fields derivable from the scalar wavefunctions of the fractional splash pulses. Hence, the fields with abnormally slow decay in the far zone can be in principle implemented as near-cycle pulses of radiofrequency, optical or acoustical waves.

### v2

@@ II Hemispherical unidirectional pulse
We calculated the Poynting vector and energy density of $U\left(\rho,z,t\right)$ according to known formulas (see, e.g., Eqs. (1)-(5) of Ref. [18]) for scalar fields, as well as the total energy. The latter is finite. The former two have normal asymptotics in the sense of finite values of the limit of the type in Eq. (2) as they have to be since both consist of derivatives of $U\left(\rho,z,t\right)$ which have normal asymptotics.
Despite the unidirectionality of $U\left(\rho,z,t\right)$—integration as a linear operation preserves the unidirectional spectrum of $u\left(\rho,z,t\right)$— in certain spatial regions the energy flux turned out to be directed opposite to the azis $z$, and at some points the velocity of such an energy backflow even reached almost the value $-c$.
@@ III Fractional splash pulse
Eqs. (15)-(17) state that if $t\rightarrow+\infty$ in other direction than along the positive axis $z$, the pulse exhibits normal asymptotics. Even if one modifies Eq. (17) by the replacements $\rho=ct\,\sin\alpha+\Delta,\,z=ct\,\cos\alpha+\Delta$, where $\alpha\approx 0$, in order to direct the limit taking almost along the axis $z$, the expression of the limit remains finite. This is explained by the relative narrowing of the transverse width of the pulse with the propagation distance $z=ct$ and the remarkable feature that only the top of the pulse decays abnormally with distance. Specifically, if we calculate in 3D the solid angle $\Omega=\pi(HWHM)^{2}/(ct)^{2}$ formed by a circular area with radius equal to transverse half width at half maximum (HWHM) of the pulse peak (of the real part) in the expanding sphere of radius $ct$, then we get the following numerical results. When $ct=10$ as in Fig. 3, then $\Omega\approx 2\,sr$; when $ct=100$, then $\Omega\approx 0.2\,$; when $ct=1000$, then $\Omega\approx 0.02\,sr$, and when $ct=10000$, then $\Omega\approx 0.002\,sr$. Hence, although the value of the HWHM increases with propagation, the angular half-width of the peak of the pulse decreases 10-fold if the propagation distance increases 10-fold and at infinity turns to zero around the axis $z$.
@@ V Discussion and study of physical feasibility
The following results were obtained : for all possible choices of $\vec{m}$, excluding $(0,0,1)$, the $x$-components and $y$-components of electric and magnetic fields both have abnormal asymptotics when $z\rightarrow\infty$ while the $z$-components do not. Hence, $E_{x}$, $E_{y}$, $B_{x}$, and $B_{y}$ behave similarly to $f\left(\rho,z,t\right)$, see Eqs. (14), (15), (28), and (29). In other words, when propagating in the positive direction of the axis $z$ they decay as $\sim 1/(z=ct)^{3/4}$, i.e., slower than the normal decay $\sim 1/(z=ct)$. If instead of the power $3/4$ another value between $1/2$ and $1$ is used in the definition of the splash pulse, the law of the decay changes correspondingly.
Fig. 7 illustrating the behavior of $E_{x}$ shows that in comparison with the scalar-valued wave function in Fig. 3, the pulse’s maximum has become more salient, and the solid angle formed by a circle with radius equal to HWHM of the pulse is more than six times smaller than that of the scalar pulse resulting in $\Omega\approx 0.0003\,sr$ if $ct=10000$ (see Section III). This is due to the circumstance that according to Eq. (30) the electromagnetic fields are expressed from $f\left(x,y,z,t\right)$ through double temporal and spatial derivatives which take their highest values in the region of the pulse maximum. It is interesting to note that only $\partial^{2}_{z,z}$ and $\partial^{2}_{t,z}$ lead to terms with abnormal asymptotics.
@@
The fact that the fractional splash pulse is not unidirectional does not imply that its physical feasibility is impossible. Its backward-propagating components can be suppressed through appropriate adjustments of the parameters $a_{1}$ and $a_{2}$. A practical approach to address this issue is to generate unidirectional finite-aperture approximations of the pulse, similar to what was done in the experimental realization of the focus wave mode [28, 29].
As to the primitives in Eqs. (3) and (9), the electromagnetic and acoustic physical fields derived from them possess normal asymptotics. Therefore, these primitives are mainly of theoretical interest.
@@ VI Conclusion
The solutions to the scalar wave equation with unusual asymptotic behavior in the far zone considered in this article are clearly of interest in mathematical physics. It is remarkable that there are several solutions that decay with propagation distance more slowly than the common inverse proportional law, yet still possess finite energy. Especially interesting is the family of fractional splash pulses. Indeed, it can be easily verified that if one changes in Eq. (13) the power $3/4$ with any other value between $1/2$ and 1, the pulse has a slower than normal decay and, at the same time, its wave function is square-integrable and has finite total energy. These characteristics are passed on to physical electromagnetic as well as to acoustic fields derivable from the scalar wavefunctions of the fractional splash pulses. Hence, the fields with abnormally slow decay in the far zone can in principle be implemented as near-cycle pulses of radiofrequency, optical, or acoustical waves.